OpenStudy (anonymous):

Find the 7th term of a geometric sequence with t1=5 and r = -2.

jimthompson5910 (jim_thompson5910):

t1 = 5 is the first term ie, 5 is the first term

jimthompson5910 (jim_thompson5910):

r = -2 is the common ratio it tells you how to get the next term

jimthompson5910 (jim_thompson5910):

t2 is the next term you multiply t1 by r to get t2 t2 = r*t1 t2 = -2*5 t2 = -10 So the second term is -10

jimthompson5910 (jim_thompson5910):

what is t3?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

@jim_thompson5910 am i right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

how about t4?

OpenStudy (anonymous):

-40?

jimthompson5910 (jim_thompson5910):

good, you've got the hang of this

jimthompson5910 (jim_thompson5910):

you keep going until you reach t7

OpenStudy (anonymous):

i get it now :) thank you!

jimthompson5910 (jim_thompson5910):

you're welcome